A mathematics placement test is given to all entering freshmen at Eastern

University. n = 58,x ̄= 4,y ̄ = 12, ∑ni=1 xi2 =232, ∑ni=1 xiyi =318. , .Fit a simple

linear regression model between x and y.


1
Expert's answer
2021-02-01T06:35:35-0500

"n = 58 \\

\bar{x} = 4 \\

\bar{y} = 12 \\

x_i = X_i - \bar{x} \\

y_i = Y_i - \bar{y} \\

y = β_1 + β_2x \\

β_2 = \frac{\sum x_iy_i}{\sum x_i^2} \\

= \frac{318}{232} = 1.37 \\

β_1 = y - β_2x \\

= 12 - (1.37 \times 4) \\

= 6.52"

The simple linear regression model between x and y:

y = 6.52 + 1.37x